# library.bib

### v1

@@ III-E Effect of Control Barrier Functions
We further illustrate the effect of CBFs in a lateral push experiment, where the robot recovers from a leftward force applied to the torso while the left leg is in stance and the right leg is in swing. This scenario is challenging because the right swing leg must step as close as possible to the left stance leg without colliding to avoid falling. Fig. 7 shows the distance constraint between the feet for different strategies: ADMM without CBFs, ADMM with CBFs, and the relaxed-barrier penalty method from [grandiaPerceptiveLocomotionNonlinear2023] for a disturbance applied at 0.35 s. For ADMM without CBFs, the distance constraint is violated by 3.7 cm at 0.45 s. However, sufficient SQP iterations allow the robot to recover with feasible narrow stepping that minimizes lateral torso velocity after 2 s. The penalty method achieves lower constraint violation (1.6 cm), but couples cost minimization and constraint satisfaction, leading to suboptimal wider steps after 1.5 s. Our approach, combining which combines ADMM and CBFs, maintains strict constraint satisfaction without altering the optimal narrow stepping behavior after recovering recovery.
@@ IV Hardware Experiments
This section validates our whole-body NMPC approach on hardware. The hardware results are best captured in the supplementary video accessible on YouTube [^1] https://youtu.be/Xmi7AFGhlFc. The NMPC allows the MIT Humanoid to holistically reason about all degrees of freedom while respecting joint limits and self-collision constraints with sufficient accuracy. This enables the robot to coordinate its arm and leg motions to dynamically balance when walking, stopping, and recovering from significant disturbances (Fig. 1). In practice, joint position limits are always respected, allowing the robot to walk stably with its knees fully extended. Combined with the sphere-based collision model from Fig. 2b and CBFs, the robot can reliably plan complex crossed-leg motions that enhance stability in challenging push recovery, abrupt halting, or fast turning scenarios.
As long as they run sufficiently fast, the different NMPC formulations transfer seamlessly to hardware. Table I shows the solve times of four NMPC formulations extensively tested on hardware with 20 ADMM iterations. The full collision model (Fig. 2b) only increases solve times by 9%, and the CBFs only increase solve times by 2.7%. Note that the standard deviations are low due to the fixed number of iterations. The formulation with $dt=50$ ms is the fastest one that remains stable on hardware and can walk at up to 0.8 m/s at a 200 Hz update rate. The formulation with $dt=25$ ms is the most successful controller (Fig. 4) that is still fast enough for online only (90 Hz). It can walk at speeds of up to 1.1 m/s on hardware and is much more resistant to push recoveries and rough terrain.
@@ V Conclusion and Outlook
This paper presented an implementation of whole-body NMPC that plans trajectories for all the degrees of freedom of the MIT Humanoid in real time. Our simulation results suggest that real robots may not benefit from high-accuracy solutions due to unavoidable inertial modeling errors, dynamics discretization errors and computation delay. Instead of seeking high accuracy, we leverage ADMM’s ability to provide reliable, low-accuracy solutions at a high frequency and show their adequacy through extensive hardware experiments. Furthermore, we use CBFs at the first timestep of the NMPC to enhance the closed-loop feasibility of self-collision constraints with minimal added computational burden. Our approach enables the MIT Humanoid to reliably walk and respond to significant disturbances with holistic whole-body reasoning while respecting contact constraints, joint limits, and self-collisions. This allows the robot to execute arm and crossed-leg motions that enhance stability.
Despite its empirical reliability, our approach cannot guarantee stability or feasibility of the solution. Additionally, the proposed approach follows a fixed contact sequence, making the controller vulnerable to larger pushes when applied at the beginning of a swing phase. Finally, the most successful controllers in simulation are still too slow for hardware deployment. Our solver can be accelerated by state-of-the-art advancements such as analytical derivatives [carpentierAnalyticalDerivativesRigid2018], parallelization [katayamaEfficientSolutionMethod2021], and structure-exploiting Riccati factorizations [jordanaStagewiseImplementationsSequential].
VI Acknowledgments
We thank Andrew SaLoutos, Hongmin Kim and David Nguyen for continuously robustifying the hardware.

### v2

@@ III-E Effect of Control Barrier Functions
We further illustrate the effect of CBFs in a lateral push experiment, where the robot recovers from a leftward force applied to the torso while the left leg is in stance and the right leg is in swing. This scenario is challenging because the right swing leg must step as close as possible to the left stance leg without colliding to avoid falling. Fig. 7 shows the distance constraint between the feet for different strategies: ADMM without CBFs, ADMM with CBFs, and the relaxed-barrier penalty method from [grandiaPerceptiveLocomotionNonlinear2023] for a disturbance applied at 0.35 s. For ADMM without CBFs, the distance constraint is violated by 3.7 cm at 0.45 s. However, sufficient SQP iterations allow the robot to recover with feasible narrow stepping that minimizes lateral torso velocity after 2 s. The penalty method achieves lower constraint violation (1.6 cm), but couples cost minimization and constraint satisfaction, leading to suboptimal wider steps after 1.5 s. Our approach, which combines ADMM and CBFs, maintains strict constraint satisfaction without altering the optimal narrow stepping behavior after recovery.
@@ IV Hardware Experiments
The hardware results are best captured in the supplementary video linked in Fig. 1. The NMPC allows the MIT Humanoid to holistically reason about all degrees of freedom while respecting joint limits and self-collision constraints with sufficient accuracy. This enables the robot to coordinate its arm and leg motions to dynamically balance when walking, stopping, and recovering from significant disturbances (Fig. 1). In practice, joint position limits are always respected, allowing the robot to walk stably with its knees fully extended. With the sphere-based collision model from Fig. 2b and CBFs, the robot reliably plans complex crossed-leg motions that enhance stability in challenging push recovery, abrupt halting, or fast turning scenarios.
The different NMPC formulations transfer seamlessly to hardware. Table I shows the solve times of four NMPC formulations extensively tested on hardware with 20 ADMM iterations. The full collision model (Fig. 2b) and the CBFs only increases solve times by 9% and 2.7%, respectively. Note that the standard deviations are low due to the fixed number of iterations. The formulation with $dt=50$ ms is the fastest one that remains stable on hardware and can walk at up to 0.8 m/s at a 200 Hz update rate. The formulation with $dt=25$ ms is the most successful controller (Fig. 4) that is still fast enough for online use (90 Hz). It can walk at speeds of up to 1.1 m/s on hardware and is much more resistant to push recoveries and rough terrain.
@@ V Conclusion and Outlook
This paper presented an implementation of whole-body NMPC that plans trajectories for all the degrees of freedom of the MIT Humanoid in real time. Our simulation results suggest that real robots may not benefit from high-accuracy solutions due to unavoidable inertial modeling errors, dynamics discretization errors and computation delay. Instead of seeking high accuracy, we leverage ADMM’s ability to provide reliable, low-accuracy QP subproblem solutions at a high frequency and show their adequacy through extensive hardware experiments. Furthermore, we use CBFs at the first timestep of the NMPC to enhance the closed-loop feasibility of self-collision constraints with minimal added computational burden. With our approach, the MIT Humanoid reliably walks and responds to significant disturbances with holistic whole-body reasoning while respecting contact constraints, joint limits, and self-collisions. This allows the robot to execute arm and crossed-leg motions that enhance stability.
Despite its reliability, our approach cannot guarantee stability or feasibility of the solution. Additionally, the controller follows a fixed contact sequence. This makes the robot vulnerable to large pushes applied at the beginning of a swing phase, where the time until the next foot touchdown is maximal. Finally, the most successful controllers in simulation are still too slow for hardware deployment. Our solver can be accelerated by state-of-the-art advancements such as analytical derivatives [carpentierAnalyticalDerivativesRigid2018], parallelization [katayamaEfficientSolutionMethod2021], and Riccati factorizations [jordanaStagewiseImplementationsSequential2023].
Acknowledgments
We thank Andrew SaLoutos, Hongmin Kim and David Nguyen for continuously robustifying the MIT Humanoid hardware.